# Formation of clusters and coarsening in weakly interacting diffusions

### v1

@@ 4.3. Application to the mean-field equation and dynamic metastability
$h_{a,b}(x)$ $=\frac{\int_{x}^{b}e^{V_{\rm eff}(y)}\mathop{}\!\mathup{d}y}{\int_{a}^{b}e^{V_{\rm eff}(y)}\mathop{}\!\mathup{d}y}\qquad\text{for all }x\in[a,b],$ (4.25)
so that as above we can approximate the equilibrium potential for all $x\in(X^{(j)}-\ell s_{w},X^{(j)}+\ell s_{w})$ by
$h_{a,b}(x)\approx p^{(j)}:=\frac{X^{(j+1)}-X^{(j)}-2s_{w}\ell}{X^{(j+1)}-X^{(j-1)}-4s_{w}\ell}.$ (4.26)
With this, we describe the mass exchange as follows: A particle $X^{i}_{t}$ of cluster $j$ waits for an exponentially distributed time with expectation $\mathbb{E}\left[\tau_{a,b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]$ inside the cluster $j$. When it leaves the cluster after this random time, it joins the cluster j1j-1 with probability $p^{(j)}$ and the cluster $j+1$ with probability $1-p^{(j)}$. This gives the following waiting times for jumps of the particle $X^{i}_{t}$ inside cluster $j$ to the cluster j1j-1 and $j+1$, respectively:
@@
$T_{\rm right}^{(j)}$ $=\frac{\mathbb{E}\left[\tau_{a,b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]}{1-p^{(j)}}\approx\left(d_{\rm dir}\left(X^{(j+1)},X^{(j)}\right)-2s_{w}\ell\right)\sqrt{\frac{2\pi\ell}{e\gamma w^{\prime\prime}(0)m^{(j)}}}e^{\gamma\ell\Delta m^{(j)}}$ (4.28)
@@ 4.6. Modeling the mass exchange via jumps of particles
Based on the waiting times for particles from one cluster to join adjacent clusters, we can write down a continuous-time stochastic Markov process for the mass exchange of the clusters which features jumps of particles from one cluster to another. Assuming again that the particles are arranged in $K(t)\in\mathbb{N}^{+}$ clusters at time $t\geq 0$ with masses $m^{(j)}=M^{(j)}/N$ and cluster centers $X^{(j)}$ where $j=1,\dots,K(t)$, we model the mass exchange of these clusters by a continuous time Markov chain with state space
@@
Figure $4.$ Free energy landscape for a two-cluster state (3.40) in dependence of the two masses masses $m_{1}\in[0,1]$, $m_{2}=1-m_{1}$ with distance of cluster centers $d:=X^{(1)}-X^{(2)}\in[0,1]$. For fixed $d$, the free energy is maximized at $m_{1}=m_{2}=1/2$, which is however an unstable state. The free energy was computed by discretizing the torus with $2^{9}$ grid points. Parameters: $\gamma=150,\ell=0.1$, $w$ as in (1.4).
@@
A multi-cluster state with $K$ clusters of the same masses $\frac{1}{K}$ and at equidistant positions on the torus corresponds to a stationary state of the ODE (4.8b). Indeed, we can prove the existence of stationary states of (1.6) of this form: By Section 3.1, for $\ell\ll 1$ and $\gamma\gg 1$, there exists a non-trivial global minimizer $q_{\gamma,\ell}$ of the free energy. Moreover, $q_{\gamma,\ell}$ must be a single-cluster state by Theorem 3.3. We show now how to construct multi-cluster stationary states from $q_{\gamma,\ell}$.
We view $q_{\gamma,\ell}$ as $1$-periodic function $\mathbb{R}\to\mathbb{R}$. Then, for every $K\in\mathbb{N}^{+}$ such that $s_{w}\ell<1/2$, the state $\tilde{q}(x):=q_{\gamma,\ell}(Kx)$ is a fixpoint of (3.2) for $W_{\tilde{\gamma},\tilde{\ell}}$ where $\tilde{\gamma}=K^{2}\gamma$ and $\tilde{\ell}=\ell/K$, since
@@ 5. Discussion and outlook
In this paper, the dynamics of weakly interacting diffusions in an attractive localized interaction potential was studied. It was shown, through a combination of heuristics, rigorous analysis, and careful numerical experiments, that the system exhibits very rich dynamical behavior that the mean-field description cannot capture. In particular, we identified several timescales and dynamical phenomena: formation of clusters, merging/coalescence, mass transfer between clusters, and microscopic reversibility, which we summarize as follows:
@@
  • Microscopic reversibility: The finite particle system (1.1) is ergodic due to the presence of Brownian motion for each particle. However, the presence of a discontinuous phase transition (see Corollary 3.2) implies that the free energy (1.7) has a critical point. Since the clustered state is the global minimizer, its energy barrier $\Delta_{\mathcal{F}}>0$ to this critical point determines the timescale of microscopic reversibility $t_{\rm reversibility}\asymp e^{N\Delta_{\mathcal{F}}}$ (see [GS20, Section 5]).
@@ Appendix A Convergence towards the Modified Massive Arratia flow
As explained above, in the regime $\gamma\gg 1$ and $\ell\ll 1$, once clusters have arisen, the system (1.1) behaves essentially as a system of coalescing heavy Brownian motions. It is therefore natural to expect that appropriate rescalings of (1.1) converge in the large particle limit to the Massive Modified Arratia Flow [KR19], which is an random element $\{y(u,t),\ u\in[0,1],\ t\in[0,T]\}$ in the Skorohod space $D([0,1],C[0,T])$ characterized by the following set of axioms.

### v2

@@ 4.3. Application to the mean-field equation and dynamic metastability
$h_{a,b}(x)$ $=\frac{\int_{x}^{b}e^{V_{\rm eff}(y)}\mathop{}\!\mathup{d}y}{\int_{a}^{b}e^{V_{\rm eff}(y)}\mathop{}\!\mathup{d}y}\qquad\text{for all }x\in[a,b].$ (4.20)
By carefully examining the shape of the effective potential $V_{\rm eff}$, Laplace’s method yields the following result for the limit $\gamma\to\infty$.
Proposition 4.1.
In addition to 1.1, suppose that $w$ is $C^{3}$ on $r_{w},r_{w}]$ for some $r_{w}\in(0,s_{w}]$ and that there exists some $\delta_{w}\in(0,s_{w})$ and $\tau>0$ such that
$w(z)\leq\tau\cdot(z-s_{w})\quad\text{for all }z\in[\delta_{w},s_{w}].$ (4.21)
Let $K\geq 3$ be a natural number, $\gamma>0$, $\ell>0$ and fix $j\in\{2,\dots,K-1\}$. Let the cluster positions $X^{(1)}<\dots<X^{(K)}\in\mathbb{R}$ and masses $m^{(1)},\dots,m^{(K)}\in(0,1)$ be given such that $\sum_{j=1}^{K}m^{(j)}=1$ and such that $X^{(k+1)}-X^{(k)}>2s_{w}\ell$ for all $k=1,\dots,K-1$. Let $\alpha\in(0,s_{w})$ such that
$\frac{\lvert w(\alpha)\rvert}{\Delta}<\min\left\{\frac{m^{(j)}}{m^{(j-1)}},\frac{m^{(j)}}{m^{(j+1)}}\right\}$ (4.22)
and define [^3] The condition (4.22) ensures that $a$ is sufficiently close to the right boundary of the potential well around $X^{(j-1)}$ and that $b$ is sufficiently close to the left boundary of the potential well around $X^{(j+1)}$. $a:=X^{(j-1)}+\alpha\ell$ and $b=X^{(j+1)}-\alpha\ell$.
Define the Gaussian mixture
$\rho({\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}}):=\sum_{k=1}^{K}m^{(k)}g({\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}};X^{(k)},\sigma_{k}^{2}),\quad\text{where }\sigma_{k}^{2}:=\frac{\ell}{\gamma w^{\prime\prime}(0)m^{(k)}},$ (4.23)
and define the effective potential
$V_{\rm eff}(x):=\left(\gamma\ell w\left(\frac{{\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}}}{\ell}\right)\ast\rho\right)(x).$ (4.24)
Consider a diffusive particle $X^{i}_{t}$ on the real line moving according to ${\rm d}X^{i}_{t}=-V^{\prime}_{\rm eff}(X^{i}_{t}){\rm d}t+\sqrt{2}{\rm d}B^{i}_{t}$ in the effective potential (4.24) which is formed by the multi-cluster state $\rho$. Consider the hitting time
$\tau_{a,b}:=\inf\{t\geq 0:X^{i}_{t}\in(-\infty,a]\cup[b,\infty)\}.$ (4.25)
Fix $x_{0}\in(X^{(j)}-s_{w}\ell,X^{(j)}+s_{w}\ell)$. Then, we have the asymptotic equivalence
  E[ τ_a,b  — X^i_0 = x_0]
(X(j+1)- X(j)-2sw) (X(j)- X(j-1)-2sw) (X(j+1)- X(j-1)- 2sw) γw”(0) m(j) e^γΔm^(j) - 12
as $\gamma\to\infty$, which is an Eyring–Kramers formula for the expected hitting time of the j1j-1-th and the $j+1$-th cluster when starting inside the $j$-th cluster. Moreover, we obtain for the equilibrium potential
$\mathbb{P}\left[\tau_{a}<\tau_{b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]\sim p^{(j)}:=\frac{X^{(j+1)}-X^{(j)}-2s_{w}\ell}{X^{(j+1)}-X^{(j-1)}-4s_{w}\ell}\quad\text{as }\gamma\to\infty.$ (4.26)
We will prove Proposition 4.1 in Appendix B. The central idea is to use a rigourous version of (4.5) to approximate $V_{\rm eff}$ around its minima $X^{(1)},\dots,X^{(K)}$. Once a particle has left a valley formed by one of the clusters, the time it takes to reach a neighbouring cluster is determined by the widths of the potential ridges between two clusters. As we will show, these widths are given by $X^{(j+1)}-X^{(j)}-2s_{w}\ell$ in the limit $\gamma\to\infty$ for the ridge between cluster $j$ and cluster $j+1$. This explains the prefactors in (LABEL:eq:clusters-masses:Eyring-Kramers).
With (LABEL:eq:clusters-masses:Eyring-Kramers) and (4.18), we can describe the mass exchange as follows: A particle $X^{i}_{t}$ of cluster $j$ waits for an exponentially distributed time with expectation $\mathbb{E}\left[\tau_{a,b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]$ inside the cluster $j$. When it leaves the cluster after this random time, it joins the cluster j1j-1 with probability $p^{(j)}$ and the cluster $j+1$ with probability $1-p^{(j)}$. This gives the following waiting times for jumps of the particle $X^{i}_{t}$ inside cluster $j$ to the cluster j1j-1 and $j+1$, respectively:
@@
$T_{\rm right}^{(j)}$ $=\frac{\mathbb{E}\left[\tau_{a,b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]}{1-p^{(j)}}\approx\left(d_{\rm dir}\left(X^{(j)},X^{(j+1)}\right)-2s_{w}\ell\right)\sqrt{\frac{2\pi\ell}{e\gamma w^{\prime\prime}(0)m^{(j)}}}e^{\gamma\ell\Delta m^{(j)}}$ (4.28)
@@ 4.6. Modeling the mass exchange via jumps of particles
Based on the waiting times for particles from one cluster to join adjacent clusters, we can write down a continuous-time stochastic Markov process for the mass exchange of the clusters which features jumps of particles from one cluster to another. Assuming again that the particles are arranged in $K(t)\in\mathbb{N}^{+}$ clusters at time $t\geq 0$ with masses $m^{(j)}=M^{(j)}/N$ and cluster centres $X^{(j)}$ where $j=1,\dots,K(t)$, we model the mass exchange of these clusters by a continuous time Markov chain with state space
@@
Figure $4.$ Free energy landscape for a two-cluster state (3.61) in dependence of the two masses masses $m_{1}\in[0,1]$, $m_{2}=1-m_{1}$ with distance of cluster centres $d:=X^{(1)}-X^{(2)}\in[0,1]$. For fixed $d$, the free energy is maximized at $m_{1}=m_{2}=1/2$, which is however an unstable state. The free energy was computed by discretizing the torus with $2^{9}$ grid points. Parameters: $\gamma=150,\ell=0.1$, $w$ as in (1.4).
@@
A multi-cluster state with $K$ clusters of the same masses $\frac{1}{K}$ and at equidistant positions on the torus corresponds to a stationary state of the ODE (4.7b). Indeed, we can prove the existence of stationary states of (1.6) of this form: By Section 3.1, for $\ell\ll 1$ and $\gamma\gg 1$, there exists a non-trivial global minimiser $q_{\gamma,\ell}$ of the free energy. Moreover, $q_{\gamma,\ell}$ must be a single-cluster state by Theorem 3.4. We show now how to construct multi-cluster stationary states from $q_{\gamma,\ell}$.
We view $q_{\gamma,\ell}$ as $1$-periodic function $\mathbb{R}\to\mathbb{R}$. Then, for every $K\in\mathbb{N}^{+}$ and assuming that $s_{w}\ell<1/2$, the state $\tilde{q}(x):=q_{\gamma,\ell}(Kx)$ is a fixpoint of (3.2) for $W_{\tilde{\gamma},\tilde{\ell}}$ where $\tilde{\gamma}=K^{2}\gamma$ and $\tilde{\ell}=\ell/K$, since
@@ 5. Discussion and outlook
In this paper, the dynamics of weakly interacting diffusions in an attractive localized interaction potential was studied. It was shown, through a combination of heuristics, rigorous analysis, and careful numerical experiments, that the system exhibits very rich dynamical behaviour that the mean-field description cannot fully capture. In particular, we identified several timescales and dynamical phenomena: formation of clusters, merging/coalescence, mass transfer between clusters, and microscopic reversibility, which we summarize as follows:
@@
  • Microscopic reversibility: The finite particle system (1.1) is ergodic due to the presence of Brownian motion for each particle. However, the presence of a discontinuous phase transition (see Proposition 3.2) implies that the free energy (1.7) has a critical point. Since the clustered state is the global minimiser, its energy barrier $\Delta_{\mathcal{F}}>0$ to this critical point determines the timescale of microscopic reversibility $t_{\rm reversibility}\asymp e^{N\Delta_{\mathcal{F}}}$ (see [GS20, Section 5]).
@@ Appendix A Convergence towards the Modified Massive Arratia flow
As explained above, in the regime $\gamma\gg 1$ and $\ell\ll 1$, once clusters have arisen, the system (1.1) behaves essentially as a system of coalescing heavy Brownian motions. It is therefore natural to expect that appropriate rescalings of (1.1) converge in the large particle limit to the Massive Modified Arratia Flow (MMA) [KR19], which is an random element $\{y(u,t),\ u\in[0,1],\ t\in[0,T]\}$ in the Skorohod space $D([0,1],C[0,T])$ characterized by the following set of axioms.